# Observing structural disorder induced interacting topological phase in an atom array

### v2

@@ Abstract
Topological phases of matter can appear in noninteracting systems, as in band topology, or interacting systems, such as in spin models, with their defining features typically robust against weak disorder. Intriguingly, disorder itself can also induce topological phases–exemplified by the Anderson topological insulator in noninteracting systems. Experimental studies on disorder induced topology have so far been limited to band topology. Here we report direct observations of structural disorder induced many-body interacting topological phase in an atom array at half-filling, whereby random offsets to tweezer locations forming a lattice implement structural disorder, causing fluctuating long-range dipolar interactions between tweezer confined single atoms. The ground state degeneracy in disordered configurations is detected and compared to a regular lattice. The induced topological phase is also vindicated by the spatially resolved atom-atom correlation functions for different forms of dimer compositions. By probing the quench dynamics of a highly excited state, we observe markedly slower decay of edge spin magnetization in comparison to the bulk spin one, consistent with the presence of topologically protected edge modes in disordered lattices. Our experiments open a new direction for studying the interplay between structural disorder and strongly interacting topological matter in Rydberg atom arrays.
@@
Topological phases of matter constitute a novel class of quantum states that extends beyond the conventional symmetry-breaking paradigm 1, 2, 3, 4. Since disorder is ubiquitous in either natural or synthetic systems, understanding its influence on topological phases is of fundamental interest 5, 6. Typically viewed as detrimental to ordered phase, disorder can in fact also induce topological phase–for example, in the celebrated topological Anderson insulator 7. This counterintuitive phenomenon has motivated experimental studies across a range of systems, from the classical ones of waveguide arrays 8, 9, 10, photonic crystals 11, 12, and acoustics 13, 14, to the quantum ones of atomic condensates in momentum lattices 15 and superconducting circuit 16. However, all realizations are concentrated so far on noninteracting particles or fully captured by band topology of single-particle states.
Beyond band topology widely studied in single-particles or noninteracting fermions, topological phases emerge in bosonic systems as well 17, 4, 18. While free fermions settle into ground state following exclusion principle, bosons typically form a condensate as the ground state. A nontrivial symmetry-protected topological (SPT) phase can nevertheless emerge for interacting bosons in the presence of strong correlations, as cultivated by hard-core bosons, i.e., in spin-1/2 quantum models 19. The classification of bosonic SPT phases involves analyzing a model’s quantum many-body ground state and is systematically carried out based on group cohomology in one-dimensional (1D) systems 20, 21, 22, 23. Among all SPT phases studied for interacting bosons, the Haldane phase stands out as a notable example of many-body topology 17 and is observed in several solid-state materials 24, 25. In synthetic matter, an SPT phase or equivalently the Haldane state, is observed in various systems: a Rydberg atom array 26, a Fermi-Hubbard ladder of atoms in optical lattices 27, a nanographene spin chain 28, 29, atom-based semiconductor quantum dots 30, and surface atoms 31. These quantum simulations 26, 27, 28, 29, 30, 31 are implemented in lattices, either without disorder or with weak disorder. In the presence of disorder, recent studies reveal when disorder breaks the protecting symmetry, an average symmetry may exist for all possible realizations of disorder 32, 33, 34, giving rise to an average SPT phase 35. In addition, theoretical studies predict disorder can even induce a bosonic SPT phase 36. However, the interesting effects of disorder on interacting SPT phase remain to be observed in experiments.
@@
This work reports the first experimental observation of a disorder induced bosonic average SPT phase. A 1D long-range SSH model for hard-core bosons (of infinite onsite interaction) is implemented with strong dipolar as well as van der Waals interactions between atoms. The initial atomic array forms a regular chain in a topologically trivial phase for either single or many particles. Structural disorder is introduced through random displacements of optical tweezers 37, 38, 39 otherwise centered on lattice sites. At the single-particle level, edge mode emerges once disorder strength exceeds a critical threshold, confirming the structural disorder induced topological state. At near half-filling with many particles, ground state degeneracy and correlation functions are observed, both substantiating the earlier prediction of an induced interacting topological phase 36. The observed quench dynamics from a highly excited state further implicate the presence of edge states by their significantly slower decay compared to atoms in the bulk. We therefore provide comprehensive evidence for the disorder induced many-body interacting topological phase, supported by measurements of the density distribution, correlation functions, microwave response, and quench dynamics.
Figure 1: Hard-core boson model with structural disorder. (A) Schematics of a dimerized 1D regular lattice with two sub-chains corresponding to a topologically trivial phase. Solid and dashed lines connecting two lattice sites represent intracell and intercell hoppings, respectively, with the line width indicating the relative strength of the hoppings. The filled dark blue circle denotes the inversion center. (B) The hopping between atoms $i$ and $j$ results from the dipolar exchange interaction that de-excites atom $i$ from $p$ to $s$ states while simultaneously excites atom $j$ from $s$ to $p$ states, or vice versa. The middle panel presents the angular ($\theta$ between ${\bm{r}}_{ij}$ and the magnetic field ${\bm{B}}$) dependence of hopping strength, measured between two atoms separated by 13 $13\text{ }\mathrm{\mu m}$ using microwave spectroscopy, and compared to the theoretical prediction shown as a solid line. At the magic angle $\theta_{m}\approx 54.7^{\circ}$ or $125.3^{\circ}$, hoppings vanish along the two sub-chains. Error bars representing the standard error of the means (SEM) are smaller than the filled circle diameter. (C) Numerically calculated topological invariant $P_{\mathrm{M}}$ (averaged over more than $150$ random configurations) for our model at half-filling as a function of the structural disorder strength $W$, illustrating a topological phase transition from a trivial state at $\overline{P}_{\mathrm{M}}=0$ to a nontrivial phase at $\overline{P}_{\mathrm{M}}=0.5$. (D) Single-shot fluorescence images of atoms arranged into a regular lattice (outlined by the dashed lines) at $W=0$ and disordered lattices at $W=0.2{d}$ and $W=0.4{d}$. As the disorder strength increases, topological edge states emerge, indicated by filled orange circles. For our experiments, ${d}=13\text{ }\mathrm{\mu m}$, $\Delta x=0.84{d}$, and $\Delta y=0.47{d}$; the resulting nearest-neighbor intracell hopping is about $0.8$ MHz for the regular lattice, compared to a numerically estimated intercell hopping of about 0.1 MHz.
@@
Hard-core boson model.—Our system is described by the following 1D staggered model involving two sub-chains $\alpha$ and $\beta$ (see Fig. 1A),
@@
with the vacuum and one-particle excitation states encoded into the Rydberg $|s\rangle=|0\rangle$ and $|p\rangle=\hat{b}^{\dagger}|0\rangle$ states, respectively 40. $\hat{b}_{j}$ ($\hat{b}_{j}^{\dagger}$) is the annihilation (creation) operator of a boson at site $j$, constrained by the hard-core condition $\hat{b}_{j}^{2}=(\hat{b}_{j}^{\dagger})^{2}=0$. $\hat{n}_{i}=\hat{b}_{i}^{\dagger}\hat{b}_{i}$ is the particle number operator. The terms in the parentheses describe hopping from dipolar exchanges between two Rydberg atoms at sites $i$ and $j$ with strength given by $J_{ij}=C_{3}(1-3\cos^{2}\theta_{ij})/(2r_{ij}^{3})$ 41, 42 which fluctuates with the random vector ${\bm{r}}_{ij}$ as structural disorder is introduced. $C_{3}$ is the dipolar interaction coefficient, and the external magnetic field ${\bm{B}}$ makes an angle $\theta_{ij}$ with ${\bm{r}}_{ij}$. $\hat{H}_{\mathrm{vdW}}=\sum_{i<j}V_{ij}^{\mathrm{vdW}}(1-\hat{n}_{i})(1-\hat{n}_{j})$ denotes van der Waals interactions between Rydberg $s$ states, with $V_{ij}^{\mathrm{vdW}}=-C_{6}/r_{ij}^{6}$ and $C_{6}$ the corresponding coefficient. The much weaker interactions between $p$ states are ignored. $V_{ij}^{\mathrm{vdW}}$ is comparable to $J_{ij}$ in our experiment, and contributes an Ising term of the form $\sum_{i<j}\sigma_{i}^{z}\sigma_{j}^{z}$ 43, 44, 45. Consequently, our model in Eq. (1) resembles the Heisenberg model with a $\sigma^{z}$ term in the spin language 46.
@@
The tweezers can create a regular lattice with each unit cell containing two sites separated by $\Delta x$ along the direction of the chain, as shown in Fig. 1D. The atomic locations in the two sub-chains labeled $\alpha$ and $\beta$ are denoted by $x_{2j-1}=j{d}$ (lattice constant ${d}$) and $x_{2j}=j{d}+\Delta x$ ($j=1,2,\dots,N$), respectively. Since the intracell hopping significantly exceeds the intercell one (e.g., $0.8$ vs $0.1$ MHz in our experiment), a topologically trivial phase results. Structural disorder introduces random displacements to the centers of tweezers from their regular lattice sites according to $x_{2j-1}\rightarrow x_{2j-1}^{\prime}=x_{2j-1}+\delta x_{j}$ and $x_{2j}\rightarrow x_{2j}^{\prime}=x_{2j}+\delta x_{j}$, i.e., each unit cell incurs a random offset $\delta x_{j}$ along the chain, which is uniformly sampled from an interval $[-W/2,W/2]$, with $W$ characterizing the disorder strength.
For a single atom excited into the $p$ state, the physics can be discussed within the single-particle subspace spanned by the basis $\gamma=\{\hat{b}_{1}^{\dagger}|0\rangle,\hat{b}_{2}^{\dagger}|0\rangle,\dots,\hat{b}_{2N}^{\dagger}|0\rangle\}$. The corresponding Hamiltonian reduces to $[H^{\mathrm{S}}]_{ij}=J_{ij}(1-\delta_{ij})-V_{i}^{\mathrm{vdW}}\delta_{ij}$ ($1\leq i,j\leq 2N$) with $V_{i}^{\mathrm{vdW}}=-C_{6}\sum_{j\neq i}1/r_{ij}^{6}$, apart from a constant. At the "magic angle" $\theta_{m}\approx 54.7^{\circ}$ or $125.3^{\circ}$, hoppings vanish along the two sub-chains, confirmed by the measured angular dependence shown in Fig. 1B. When $V_{i}^{\mathrm{vdW}}=0$, the Hamiltonian respects the sublattice symmetry, $\Pi H^{\mathrm{S}}\Pi^{-1}=-H^{\mathrm{S}}$, also known as chiral symmetry, with $\Pi=\mathrm{diag}\{(-1)^{j}\}_{j=1}^{2N}$ the sublattice symmetry operator, leading to the $\mathbb{Z}$ classification 3. The topology of the single-particle Hamiltonian can be characterized by the polarization $P_{\mathrm{S}}$ (a topological invariant) 47, which substitutes the role played by the collective Berry phase in the regular lattice.
@@
A nonzero $V_{i}^{\mathrm{vdW}}$ breaks the sublattice symmetry. Nevertheless, for our structural disorder model, all the Hamiltonians on random lattice configurations ($\mathcal{E}_{C}$) constitute an ensemble $\mathcal{E}_{H}\equiv\left\{H^{\mathrm{S}}(C):C\in\mathcal{E}_{C}\right\}$ respecting an average inversion symmetry with respect to $x_{c}$, which is marked by the dark blue filled circle in Fig. 1A. The average symmetry arises because the configuration $C=\{x_{j}^{\prime}:j=1,2,\dots,2N\}$ occurs with the same probability as its inversion partner $\mathcal{R}C=\{\mathcal{R}x_{j}^{\prime}:j=1,2,\dots,2N\}$, where $\mathcal{R}x_{j}^{\prime}=2x_{c}-x_{j}^{\prime}$ is the inverted coordinate of $x_{j}^{\prime}$. This allows us to generalize $P_{\mathrm{S}}$ to $P_{\mathrm{S}}(C)$ for the ground states of the pair of inversion Hamiltonians $H^{\mathrm{S}}(C)$ and $H^{\mathrm{S}}(\mathcal{R}C)$, and prove that $P_{\mathrm{S}}(C)$ is quantized to 00 or $0.5$ based on the average symmetry as detailed in the Supplementary Information (SI). Since $P_{\mathrm{S}}(C)$ cannot vary continuously as the Hamiltonian pair are continuously deformed, their ground states have a topological feature with $P_{\mathrm{S}}(C)=0$ for the trivial phase and $P_{\mathrm{S}}(C)=0.5$ for the nontrivial one. Accordingly, in the single-particle case, we find increasing structural disorder strength indeed induces an average SPT phase, as illustrated in Fig. 2D.
@@
At half-filling with $N$ atoms in the $p$ state, we rewrite the hard-core boson model Eq. (1) as the Heisenberg model with a $\sigma^{z}$ term, $\hat{H}=\sum_{i<j}^{2N}[(J_{ij}/2)(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{j}^{y}\sigma_{i}^{y})+(V_{ij}^{\mathrm{vdW}}/4)\sigma_{i}^{z}\sigma_{j}^{z}]+(1/4)\sum_{i}V_{i}^{\mathrm{vdW}}\sigma_{i}^{z}$. The total spin along $z$, $\sigma^{z}=\sum_{i}\sigma_{i}^{z}$, or the $U(1)$ symmetry, is preserved as $[\hat{H},\sigma^{z}]=0$, which confines the system to the fixed $\sigma^{z}$ subspaces. When $V_{i}^{\mathrm{vdW}}=0$, the Hamiltonian also respects additional symmetries: rotations about $x$ or $y$ and time-reversal 26, 36. In a regular lattice without disorder, the resulting SPT phase is topologically equivalent to the Haldane phase and features four-fold degenerate ground states 48, 26. This phase is also characterized by a $\mathbb{Z}_{2}$ topological invariant $P_{\mathrm{M}}$ 49, which is enforced by the aforementioned symmetries to be quantized for an arbitrary spin system 50, 36. These results establish the basis for demonstrating transitions from the Affleck-Kennedy-Lieb-Tasaki (AKLT) ground state to a trivial gapped ground state 50 and an SPT phase induced by structural disorder 36.
A non-vanishing $\sigma^{z}$ term breaks all the above mentioned symmetries except for the $U(1)$ symmetry. As a result of the average inversion symmetry mentioned before and detailed in the SI, the many-body topological invariant $P_{\mathrm{M}}$, analogous to the single-particle case as detailed in the SI, is generalized to $P_{\mathrm{M}}(C)$ defined as
@@
Here $|\Psi_{S}\rangle$ is the many-body ground state of the Hamiltonian $\hat{H}$ at half-filling in the lattice configuration $S$ under periodic boundary conditions, and $\hat{\mathcal{P}}_{\mathrm{M}}(S)=\prod_{j=1}^{2N}e^{-\frac{\pi i}{N{d}}x_{j}\sigma_{j}^{z}}$ is the twist operator 49. Analogously, we prove in the SI that $P_{\mathrm{M}}(C)$ is quantized to 00 or $0.5$, and thus serves as a well-defined topological invariant. Figure 1C illustrates that averaged over many independent realizations, a sharp transition with increasing structural disorder strength indeed appears, indicating that a many-body average SPT phase is induced.
[…]
@@
Figure 2: Observation of a structural disorder induced topological phase transition at the single-particle level. (A) Single-particle spectra for regular and disordered lattices ($W=0.4{d}$). In the regular lattice, no edge state near zero energy is found, while in the disordered lattice, two-fold statistically degenerate edge states emerge (dark red lines), as shown by the local density of states (LDOS). Microwaves with a Rabi frequency $\Omega_{\mathrm{\mu w}}$ and detuning $\Delta_{\mathrm{\mu w}}$ probe the edge states as shown in A. The energy shifts for edge states and bulk states are attributed to van der Waals interactions (see SI). (B) The dependence of measured site occupancy on the detuning $\Delta_{\mathrm{\mu w}}$ for a regular lattice without edge states (left) and for disordered lattices with edge states at $W=0.4{d}$ (right). (C) The measured dependence of site occupancy at zero detuning on structural disorder strength, showing a clear increase at two edges. (D) Contrast between the excitation of edge sites and bulk sites as a function of disorder strength from the same data in (C). The gray solid line represents the numerical result incorporating various experimental errors. Error bars are SEMs averaged over $15$ random configurations. Inset: numerically computed polarization $\overline{P}_{\mathrm{S}}$ for the Hamiltonian $H^{\mathrm{S}}$ with $N=7$ (dashed line) and $N=40$ (solid line) as functions of disorder strength, averaged over $1000$ realizations. Experimental results for disordered lattices are averaged over $15$ random configurations.
@@
Using microwave spectroscopy, we determine the presence or absence of an excitation at the edge sites, thus affirmatively demonstrating the structural disorder induced topology at the single-particle level. After initializing all atoms into the $s$ state, a weak microwave field is applied, coupling it to the $p$ state in a process described by the Hamiltonian $\hat{H}_{\mathrm{\mu w}}=\sum_{i=1}^{2N}[(\hbar\Omega_{\mathrm{\mu w}}/2)(\hat{b}_{i}^{\dagger}+\hat{b}_{i})-\hbar\Delta_{\mathrm{\mu w}}\hat{b}_{i}^{\dagger}\hat{b}_{i}]$ with the Rabi frequency $\Omega_{\mathrm{\mu w}}/(2\pi)\approx 0.2$ MHz and the microwave detuning $\Delta_{\mathrm{\mu w}}$. An edge excitation will be created when the detuning matches the resonance as illustrated in Fig. 2A after 2.5 $2.5\text{ }\mathrm{\mu s}$. Figure 2B presents the $p$ state occupancy in the end of excitation, revealing pronounced peaks near zero detuning at the edges of disordered lattices with $W=0.4{d}$. In contrast, the same edge occupancy in the regular lattice is found to resemble closely that of the bulk. The van der Waals interactions are found to slightly modify the energy at the edges but significantly shift the bulk states, although they do not break the topological nature of the system (see SI). Our observations thus confirm the presence of edge states near zero energy in the disordered lattice as well as their absence in the regular lattice.

### v3

# Observation of average topological phase in disordered Rydberg atom array

@@ Abstract
Topological phases have been extensively studied over the past two decades, primarily in quantum pure states, where they are protected by exact symmetries. Recently, numerous studies have theoretically demonstrated the existence of average symmetry-protected topological (SPT) phases in mixed quantum states, which naturally arise in real systems due to decoherence or disorder. Despite extensive experimental observations of exact SPT phases in various systems, ranging from solid-state materials to synthetic matters, average SPT phases are yet to be observed until this work. Here we report direct observations of disorder-induced many-body interacting average SPT phase in an atom array at half-filling, whereby random offsets to tweezer locations forming a lattice implement structural disorder, resulting in fluctuating long-range dipolar interactions between tweezer confined single atoms. The induced topological phase is vindicated by the spatially resolved atom-atom correlation functions for different forms of dimer compositions. The ground state degeneracy in disordered configurations is detected and compared to the regular lattice without disorder. By probing the quench dynamics of a highly excited state, we observe markedly slower decay of edge spin magnetization in comparison to the bulk spin, consistent with the presence of topologically protected edge modes in disordered lattices.
The classification of topological phases, such as topological insulators 1, 2 and the Haldane phase 3, relies critically on symmetries in pure quantum states 4, 5, 6, 7, 8, 9. Yet, real systems–from solid-state materials with inherent disorder to noisy intermediate-scale quantum (NISQ) simulators 10 with decoherence–host mixed quantum states that evade the pure-state framework. This has spurred significant recent interest in defining and understanding topological phases in mixed states 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24. A key concept developed concerns average (or weak) symmetry, which ensures that a mixed state’s density matrix is invariant under a symmetry’s action even if its constituent pure states are not 13, 14, 16, 19, in stark contrast to exact symmetry. Such symmetry can give rise to average symmetry-protected topological (SPT) phases in both noninteracting systems 11, 12, 13, 14, 15 and quantum spin models 16, 17, 18, 19, 21, 23, 24. While the exact SPT phases in both noninteracting systems and quantum spin models have been experimentally observed in various solid-state 25, 26, 27, 28, 29 and synthetic platforms 30, 31, 32, 33, 34, 35, 36, 37, an average SPT phase remains experimentally elusive despite extensive theoretical studies.
@@
Contrary to the expectation that disorder is harmful to topological phases, earlier studies have shown that it can actually induce exact SPT phases from a trivial one in both noninteracting systems (e.g., the well-known topological Anderson insulator 38) and quantum spin models 39, provided that disorder does not break the protecting symmetry. This counterintuitive phenomenon has motivated experimental studies across a range of systems, from the classical ones of waveguide arrays 40, 41, 42, photonic crystals 43, 44, and acoustics 45, 46, to the quantum ones of atomic condensates in momentum lattices 47 and superconducting circuit 48. However, all realizations are concentrated so far on exact SPT phases for noninteracting particles or are fully captured by band topology of single-particle states.
Figure 1: Hard-core boson model with structural disorder. a, Schematics of a dimerized 1D regular lattice with two sub-chains corresponding to a topologically trivial phase. Structural disorder is introduced by randomly displacing each unit cell from their regular lattice sites along the sub-chain direction, as illustrated in the inset with $d=13\text{ }\mu\text{m}$ in our experiments. The light red tubes denote the 1013 nm addressing laser beams used to prepare a Néel state, and $B$ represents the magnetic field. b, Interaction between Rydberg atoms: the flip-flop hopings between atoms $i$ and $j$ result from the dipolar exchange interaction that de-excites atom $i$ from $p$ to $s$ states while simultaneously excites atom $j$ from $s$ to $p$ states, or vice versa. Two Rydberg $s$ atoms also experience van der Waals (vdW) interaction, which shifts the energy of the pair states ssss. c, Numerically calculated topological invariant $P_{\mathrm{M}}$ (averaged over more than $150$ random configurations) for our model at half-filling as a function of the structural disorder strength $W$, illustrating a topological phase transition from a trivial state at $\overline{P}_{\mathrm{M}}=0$ to a nontrivial phase at $\overline{P}_{\mathrm{M}}=0.5$. d, Illustration of the emergence of inversion symmetry on average over an ensemble of all disordered Hamiltonians, in contrast to the absence of exact inversion symmetry for a Hamiltonian on a single disordered configuration. This occurs because every configuration $C$ is equally probable as its inversion partner $\mathcal{R}C$, rendering the ensemble of all disordered Hamiltonians statistically symmetric under inversion.
@@
Here we theoretically predict and experimentally realize a disorder-induced bosonic average SPT phase in a Rydberg atom array. Specifically, we implement a 1D long-range Su-Schrieffer-Heeger (SSH) model for hard-core bosons (of infinite onsite interaction) with strong dipolar besides van der Waals interactions between atoms. The initial atomic array forms a regular chain in a topologically trivial phase for either single or many particles. Structural disorder is introduced through random displacements of optical tweezers 49, 50, 51 otherwise centered on regular lattice sites. At the single-particle level, edge mode emerges once disorder strength exceeds a critical threshold, confirming the structural disorder-induced topological state. At near half-filling with many particles, ground state degeneracy and correlation functions are observed, both substantiating the earlier prediction of an induced interacting topological phase 39. The observed quench dynamics from a highly excited state further implicate the presence of edge states by their significantly slower decay compared to atoms in the bulk. We therefore provide comprehensive evidence for the disorder-induced many-body interacting topological phase, supported by measurements of the density distribution, correlation functions, microwave response, and quench dynamics.
@@
Our system is described by the following 1D staggered model involving two sub-chains $\alpha$ and $\beta$ (see Fig. 1a),
@@
with the vacuum and one-particle excitation states encoded into the Rydberg $|s\rangle=|0\rangle$ and $|p\rangle=\hat{b}^{\dagger}|0\rangle$ states, respectively 52. $\hat{b}_{j}$ ($\hat{b}_{j}^{\dagger}$) is the annihilation (creation) operator of a boson at site $j$, constrained by the hard-core condition $\hat{b}_{j}^{2}=(\hat{b}_{j}^{\dagger})^{2}=0$. $\hat{n}_{i}=\hat{b}_{i}^{\dagger}\hat{b}_{i}$ is the particle number operator. The terms in the parentheses describe hopping from dipolar exchanges between two Rydberg atoms at sites $i$ and $j$ (see Fig. 1b) with strength given by $J_{ij}=C_{3}(1-3\cos^{2}\theta_{ij})/(2r_{ij}^{3})$ 53, 54 which fluctuates with the random vector ${\bm{r}}_{ij}$ as structural disorder is introduced. $C_{3}$ is the dipolar interaction coefficient, and the external magnetic field $\bm{B}$} makes an angle $\theta_{ij}$ with respect to ${\bm{r}}_{ij}$. $\hat{H}_{\mathrm{vdW}}=\sum_{i<j}V_{ij}^{\mathrm{vdW}}(1-\hat{n}_{i})(1-\hat{n}_{j})$ denotes van der Waals interactions between Rydberg $s$ states, with $V_{ij}^{\mathrm{vdW}}=-C_{6}/r_{ij}^{6}$ and $C_{6}$ the corresponding coefficient. The much weaker interactions between $p$ states are ignored. $V_{ij}^{\mathrm{vdW}}$ is comparable to $J_{ij}$ in our experiment, and contributes an Ising term of the form $\sum_{i<j}\sigma_{i}^{z}\sigma_{j}^{z}$ 55, 56, 57.
@@
The tweezers can create a regular lattice with each unit cell containing two sites separated by $\Delta x$ along the direction of the chain, as shown in Fig. 1a. The regular lattice atomic locations in the two sub-chains labeled $\alpha$ and $\beta$ are denoted by $x_{2j-1}=j{d}$ (lattice constant d{d}) and $x_{2j}=j{d}+\Delta x$ ($j=1,2,\dots,N$), respectively. Since the intracell hopping significantly exceeds the intercell one (e.g., $0.8$ vs $0.1$ MHz in our experiment), a topologically trivial phase results. Structural disorder introduces random displacements to the centers of tweezers from their regular lattice sites according to $x_{2j-1}\rightarrow x_{2j-1}^{\prime}=x_{2j-1}+\delta x_{j}$ and $x_{2j}\rightarrow x_{2j}^{\prime}=x_{2j}+\delta x_{j}$, i.e., each unit cell incurs a random offset $\delta x_{j}$ along the chain, which is uniformly sampled from an interval [W/2,W/2][-W/2,W/2], with $W$ characterizing the disorder strength (Fig. 1a).
@@
For a single atom excited into the $p$ state, the physics can be discussed within the single-particle subspace with the corresponding Hamiltonian reducing to $H^{\mathrm{S}}]_{ij}=J_{ij}(1-\delta_{ij})-V_{i}^{\mathrm{vdW}}\delta_{ij}$ ($1\leq i,j\leq 2N$), where $V_{i}^{\mathrm{vdW}}=-C_{6}\sum_{j\neq i}1/r_{ij}^{6}$, apart from a constant. At the "magic angle" $\theta_{m}\approx 54.7^{\circ}$ or $125.3^{\circ}$, hoppings vanish along the two sub-chains. When $V_{i}^{\mathrm{vdW}}=0$, the Hamiltonian respects the sublattice symmetry, leading to the $\mathbb{Z}$ classification 58. A nonzero $V_{i}^{\mathrm{vdW}}$ due to Van der Waals interactions breaks the sublattice symmetry. Nevertheless, for our structural disorder model, all the Hamiltonians on random lattice configurations ($\mathcal{E}_{C}$) constitute an ensemble $\mathcal{E}_{H}\equiv\left\{H^{\mathrm{S}}(C):C\in\mathcal{E}_{C}\right\}$ respecting an average inversion symmetry with respect to $x_{c}$, which is marked by the black filled circle in Fig. 1d. The average symmetry arises because the configuration $C=\{x_{j}^{\prime}:j=1,2,\dots,2N\}$ occurs with the same probability as its inversion partner $\mathcal{R}C=\{\mathcal{R}x_{j}^{\prime}:j=1,2,\dots,2N\}$, where $\mathcal{R}x_{j}^{\prime}=2x_{c}-x_{j}^{\prime}$ is the inverted coordinate of $x_{j}^{\prime}$. This allows us to generalize the polarization $P_{\mathrm{S}}$ (a topological invariant) 59 to $P_{\mathrm{S}}(C)$ for the ground states of the pair of inversion Hamiltonians $H^{\mathrm{S}}(C)$ and $H^{\mathrm{S}}(\mathcal{R}C)$, and prove that $P_{\mathrm{S}}(C)$ is quantized to 0 or $0.5$ based on the average symmetry as detailed in the Supplementary Information (SI). Since $P_{\mathrm{S}}(C)$ cannot vary continuously as the Hamiltonian pair are continuously deformed, their ground states have a topological feature with $P_{\mathrm{S}}(C)=0$ for the trivial phase and $P_{\mathrm{S}}(C)=0.5$ for the nontrivial one. Accordingly, in the single-particle case, we find increasing structural disorder strength indeed induces an average SPT phase, as illustrated in Fig. 2d.
@@
At half-filling with $N$ atoms in the $p$ state, we rewrite the hard-core boson model as the Heisenberg model with a $\sigma^{z}$ term 60, which represents a genuine interacting many-body system as it cannot be mapped to a free fermion model (see SI). The total spin along $z$, $\sigma^{z}=\sum_{i}\sigma_{i}^{z}$, or the $U(1)$ symmetry, is preserved as $\hat{H},\sigma^{z}]=0$, which confines the system to the fixed $\sigma^{z}$ subspaces. When $V_{i}^{\mathrm{vdW}}=0$, the Hamiltonian also respects additional symmetries: rotations about $x$ or $y$ and time-reversal 32, 39. In a regular lattice without disorder, the resulting SPT phase is topologically equivalent to the Haldane phase and features four-fold degenerate ground states 61, 32. This phase is also characterized by a $\mathbb{Z}_{2}$ topological invariant $P_{\mathrm{M}}$ 62, 63, 39. A non-vanishing $\sigma^{z}$ term breaks all the above mentioned symmetries except for the $U(1)$ symmetry. As a result of the average inversion symmetry mentioned before and detailed in the SI, the many-body topological invariant $P_{\mathrm{M}}$, analogous to the single-particle case as detailed in the SI, is generalized to $P_{\mathrm{M}}(C)$ defined as
@@
Here $|\Psi_{S}\rangle$ is the many-body ground state of the Hamiltonian $\hat{H}$ at half-filling in the lattice configuration $S$ under periodic boundary conditions, and $\hat{\mathcal{P}}_{\mathrm{M}}(S)=\prod_{j=1}^{2N}e^{-\frac{\pi i}{N{d}}x_{j}\sigma_{j}^{z}}$ is the twist operator 62. Analogously, we prove in the SI that $P_{\mathrm{M}}(C)$ is quantized to 0 or $0.5$, and thus serves as a well-defined topological invariant. Figure 1c illustrates that averaged over many independent realizations, a sharp transition with increasing structural disorder strength indeed appears, indicating that a many-body average SPT phase is induced.
Our experiments are carried out on a short array of fourteen 87Rb atoms (N=7N=764, comparable to earlier studies in an atom array 32 or quantum dots 36 or surface atoms 37. Each atom is initially prepared in the $s$ state (denoted by $\left|\uparrow\right\rangle\equiv\left|s\right\rangle\equiv|55S_{1/2},m_{J}=-1/2\rangle$) via two-photon stimulated Raman adiabatic passage (STIRAP) from the ground state $|g\rangle\equiv|5S_{1/2},F=2,m_{F}=-2\rangle$ through the intermediate state $|e\rangle\equiv|6P_{3/2},F=3\rangle$. Further excitation to the $p$ state ($\left|\downarrow\right\rangle\equiv|p\rangle\equiv|55P_{1/2},m_{J}=1/2\rangle$) is accomplished by a microwave field at a resonant frequency of about $E_{0}/h\sim 22.1$ GHz. A bias magnetic field of $B_{z}\approx 30$ G prevents cross-couplings with other magnetic sub-levels. Each experiment typically lasts for a few microseconds, during which the atomic spatial motion can be considered frozen. At the end of an experiment, atoms in the $s$ state are de-excited to the ground state, imaged by fluorescence spectroscopy, while atoms in the $p$ state are lost. All experimental results for disordered lattices are averaged over $15$ random configurations.
@@
Figure 2: Observation of a structural disorder-induced topological phase transition at the single-particle level. a, Single-particle spectra for regular and disordered lattices (W=0.4dW=0.4{d}). In the regular lattice, no edge state near zero energy is found, while in the disordered lattice, two-fold statistically degenerate edge states emerge (dark red lines), as shown by the local density of states (LDOS). Microwaves with a Rabi frequency $\Omega_{\mathrm{\mu w}}$ and detuning $\Delta_{\mathrm{\mu w}}$ probe the edge states. b, The dependence of measured site occupancy on the detuning $\Delta_{\mathrm{\mu w}}$ for a regular lattice without edge states (left) and for disordered lattices with edge states at W=0.4dW=0.4{d} (right). c, The measured dependence of site occupancy at zero detuning on structural disorder strength, showing a clear increase at two edges. d, Contrast between the excitation of edge sites and bulk sites as a function of disorder strength from the same data in c. The gray solid line represents the numerical result incorporating various experimental errors. Error bars are SEMs averaged over $15$ random configurations. Inset: numerically computed polarization $\overline{P}_{\mathrm{S}}$ with $N=7$ (dashed line) and $N=40$ (solid line) versus disorder strength, averaged over $1000$ realizations.
@@
Using microwave spectroscopy, we determine the presence or absence of an excitation at the edge sites, thus affirmatively demonstrating the structural disorder-induced topology at the single-particle level. After initializing all atoms into the $s$ state, a weak microwave field is applied, coupling it to the $p$ state in a process described by the Hamiltonian $\hat{H}_{\mathrm{\mu w}}=\sum_{i=1}^{2N}[(\hbar\Omega_{\mathrm{\mu w}}/2)(\hat{b}_{i}^{\dagger}+\hat{b}_{i})-\hbar\Delta_{\mathrm{\mu w}}\hat{b}_{i}^{\dagger}\hat{b}_{i}]$ with the Rabi frequency $\Omega_{\mathrm{\mu w}}/(2\pi)\approx 0.2$ MHz and the microwave detuning $\Delta_{\mathrm{\mu w}}$. An edge excitation is created when the detuning matches the resonance as illustrated in Fig. 2a after $2.5\text{ }\mathrm{\mu s}$. Figure 2b presents the $p$ state occupancy at the end of excitation, revealing pronounced peaks near zero detuning at the edges of disordered lattices with W=0.4dW=0.4{d}. In contrast, the same edge occupancy in the regular lattice is found to resemble closely that of the bulk. The van der Waals interactions are found to slightly modify the energy at the edges but significantly shift the bulk states, although they do not break the topological nature of the system (see SI). Our observations thus confirm the presence of edge states near zero energy in the disordered lattice as well as their absence in the regular lattice.